Prove that the divergence of a curl is always zero:

∇⃗⃗. (∇⃗⃗

×a)=0


1
Expert's answer
2020-05-11T17:17:30-0400

Let us recall the propertied of Levi-Civata's symbol "\epsilon_{ijk}" in three dimension,where "i,j,k \in \{1,2,3\}" .

Properties,


"\epsilon_{123}=\epsilon_{231}=\epsilon_{312}=1 \hspace{1cm} (\star)""\epsilon_{321}=\epsilon_{213}=\epsilon_{132}=-1 \hspace{1cm} (\star \star)"

"\epsilon_{ijk}=0 \hspace{1cm} (\dag)"


if at least two of "i,j,k" are equal.




If we exchange any two consecutive indices like "i \leftrightarrow j" then "\epsilon_{ijk}=-\epsilon_{jik} \:(\spades)"



The cross product of any two vector in terms of Levi-Civata's symbol is defined as bellow,


"[\overrightarrow{A} \times \overrightarrow{B}]_i =\sum_{j=1}^{3} \sum_{k=1}^{3}\epsilon_{ijk} A_j B_k \hspace{1cm} (\dag \dag)"

where,subscript "i ,j,k" of any vector denotes respectively "i^{th},j^{th},k^{th}" component of that vector.

Now, denote

"\partial_i = \frac{\partial}{\partial x^{i}}"

Let the vector "\overrightarrow{a} = a_1 \hat{e_1}+a_2 \hat{e_2}+a_3 \hat{e_3}" ,where "\hat{e_i}" is the orthogonal unit vector for all "i \in \{1,2,3\}" .

Thus,from "(\dag),(\dag \dag) ,(\star),(\star \star)" and above notation, we get ,

"\overrightarrow{\nabla} \cdot (\overrightarrow{\nabla} \times \overrightarrow{a})= \sum_{i=1}^{3} \partial_{i} (\sum_{j=1}^{3} \sum_{k=1}^{3}\epsilon_{ijk} \partial_j a_k)\\
\overrightarrow{\nabla} \cdot (\overrightarrow{\nabla} \times \overrightarrow{a})= \sum_{i=1}^{3} \sum_{j=1}^{3} \sum_{k=1}^{3}\epsilon_{ijk}\partial_i \partial_j a_k \\
\overrightarrow{\nabla} \cdot (\overrightarrow{\nabla} \times \overrightarrow{a})= \sum_{i=1}^{3} \sum_{j=1}^{3} \sum_{k=1}^{3}-\epsilon_{jik}\partial_j \partial_i a_k \hspace{1cm} (from, \spades)\\
\overrightarrow{\nabla} \cdot (\overrightarrow{\nabla} \times \overrightarrow{a})= \sum_{i=1}^{3} \sum_{j=1}^{3} \sum_{k=1}^{3}-\epsilon_{ijk}\partial_i \partial_j a_k\\
\overrightarrow{\nabla} \cdot (\overrightarrow{\nabla} \times \overrightarrow{a})=-\overrightarrow{\nabla} \cdot (\overrightarrow{\nabla} \times \overrightarrow{a})\\
 \implies \overrightarrow{\nabla} \cdot (\overrightarrow{\nabla} \times \overrightarrow{a})= 0"

Hence, we are done.